1. Two students are sick in bed with fevers. Bridgette has a temperature of 2 deg.C above normal, while Mike's is 2 deg.F above normal. Which student has the higher fever?
2.
A lady ate a whole medium size pizza storing up 120 food Calories. This energy is used to heat up 2 liters of water from a certain initial temperature to a final temperature of 80 deg.C, what was the initial temperature of the water?
Answer #1

1.Bridgette

2. ms (t2-t1)

= 2000 * 1 * (80 - t)= 120000

t = 20degrees
Answer #2

1.Bridgette has a higher fever because 1 deg C > 1deg F

C/5 = (F-32)/8

so faherhnite scale has more division than celcius scale.


2.h = m*S*(t2-t1)

120 = 2*1*(80 - t1)

80 - t1 = 60

t1 = 20 deg C
Answer #3

2 degree c is greater than 2 degree F